OpenStudy (anonymous):

Which of the following expressions is equivalent to the one shown below? 2log2 12- log2 4+log2 3 2log2 12 2log2 23 2log2 11 2log2 108 PLEASE HELP

OpenStudy (anonymous):

2log2 12-log2 4+ log2 3 = log2 12^2 - 1og2 4 + log2 3 =log2 (12x12x3/4) =1og2 108

OpenStudy (anonymous):

Are you 100% sure

OpenStudy (anonymous):

yes